OpenStudy (anonymous):

-x - y = 18 2x - y = -1 How do I solve this? If it helps any I'm currently solving systems of equations using the addition method however, when I try this I end up with x - 2y = 17 and I don't know what to do from there.

OpenStudy (skullpatrol):

First multiply one of the equations by -1 so that when you add the equations the y's will add to 0

OpenStudy (anonymous):

Alright, now its x + y = -18 2x - y = -1 then I use the addition method 3x = -19 x = -19/3

OpenStudy (skullpatrol):

$$-x - y = 18$$ $$2x - y = -1$$ $$-1(-x - y) =(-1) 18$$ $$x+y=-18$$ $$2x - y = -1$$

OpenStudy (anonymous):

was x = -19/3 correct?

OpenStudy (skullpatrol):

You will not be able to check that until you solve for y.

OpenStudy (skullpatrol):

Substitute x = -19/3 into one of your equations and solve for y.

OpenStudy (anonymous):

How do I do 2 * -19/3? the negative is confusing me.

OpenStudy (skullpatrol):